The measure of one angle of a triangle is twice the measure of a second angle. The measure of the third angle is 12 less than the sum of the other two. Find the measures of the angles of the triangle.
Answer by cleomenius(959)

A = 2x
B = x
C = x + 2x - 12
===========================================
2x + x + x + 2x -12 = 180 degrees.
6x = 12 = 180 degrees
6x =192
x = 28
============================================
A = 64 degrees
B = 32 degrees
C = 84 degrees
==========================================
64 + 32 + 84 = 180 degrees. This does check.
Cleomenius.